OpenStudy (dream277):

If x+3/3 = y+2/2 then ×/3 = So confused I'm getting ×=3/2y

OpenStudy (thephysicsman):

Well let's first re-write this

OpenStudy (dream277):

Okay

OpenStudy (thephysicsman):

\[\frac{ x+3 }{ 3 } = \frac{ y+2 }{ 2 }\]

OpenStudy (dream277):

Right

OpenStudy (thephysicsman):

Do you get what i'm doing so far? \[\frac{ x }{ 3 }+\frac{ 3 }{ 3 } = \frac{ y }{ 2 } + \frac{ 2 }{ 2 } \] \[\frac{ x }{ 3 } + 1 = \frac{ y }{ 2 } + 1\]

OpenStudy (dream277):

Ahh I see now I would of cross multiplied the way they were

OpenStudy (thephysicsman):

so then you would solve for (x/3)

OpenStudy (thephysicsman):

\[\frac{ x }{ 3 } = \frac{ y }{ 2 }-1+1\]

OpenStudy (dream277):

Is it not 3y/2

OpenStudy (dream277):

Or I'm doing something wrong still

OpenStudy (dream277):

I thought way you had it set up in was trying to solve for x so simplified both

OpenStudy (thephysicsman):

\[6*\frac{ x }{ 3 } = \frac{ y }{ 2 }*6\] 2x = 3y so now solve for x

OpenStudy (thephysicsman):

your answer is right it's just that y isn't in the denominator

OpenStudy (thephysicsman):

\[\frac{ 3 }{ 2 }y\] = x was this what you got?

OpenStudy (dream277):

Omg crazy how 1 little thing can mess you up thank you so much for help...I need tutor ha yes I figured that out lol yea not in denominator

OpenStudy (mathmale):

Please note: You are solving for the expression ×/3, in terms of y, and you do not have sufficient information from which to find a numerical value for x. Don't try to solve for x. Instead, solve for (x/3).